OpenStudy (anonymous):

The area of a rectangular wall in a classroom is 165 square feet. Its length is 4 feet longer than its width. Find the width of the wall of the classroom.

OpenStudy (jiteshmeghwal9):

Let the width be 'x'.then, length=4+x now ,\[area \space of \space rectangle =l \times w\]so,A/q\[(4+x)(x)=165\]can u do it from here????

OpenStudy (anonymous):

\[\sqrt{165}=(x+2)^2\]

OpenStudy (anonymous):

\[-2 +- \sqrt{165}=x\] but the answer is 11 and I have no idea how they got it

OpenStudy (anonymous):

this is a vocational in "completing the square"